# SPARTAN: A Sparse Transformer Learning Local Causation

### v2

@@ Abstract
Causal structures play a central role in world models that flexibly adapt to changes in the environment. While recent works motivate the benefits of discovering local causal graphs for dynamics modelling, in this work we demonstrate that accurately capturing these relationships in complex settings remains challenging for the current state-of-the-art. To remedy this shortcoming, we postulate that sparsity is a critical ingredient for the discovery of such local causal structures. To this end we present the SPARse TrANsformer World model (SPARTAN), a Transformer-based world model that learns local causal structures between entities in a scene. By applying sparsity regularisation on the attention pattern between object-factored tokens, SPARTAN identifies sparse local causal models that accurately predict future object states. Furthermore, we extend our model to capture sparse interventions with unknown targets on the dynamics of the environment. This results in a highly interpretable world model that can efficiently adapt to changes. Empirically, we evaluate SPARTAN against the current state-of-the-art in object-centric world models on observation-based environments and demonstrate that our model can learn accurate local causal graphs and achieve significantly improved few-shot adaptation to changes in the dynamics of the environment as well as robustness against removing irrelevant distractors.
@@ 1 Introduction
World Models [17] have emerged in recent years as a promising paradigm to enable a wide range of downstream tasks such as video prediction [58, 19], physical reasoning [6, 9] and model-based RL [18]. While recent advances in dynamics modelling have developed world models that are capable of performing accurate predictions over ever longer horizons in increasingly complex environments [38, 46], the ability to adapt to changes in the environment in a data efficient manner remains a significant challenge. To this end, the intersection between causality and machine learning [48, 25] offers promising opportunities for building structured models that can reason about changes in the environment. In particular, causal graphical models [41] afford the notion of interventions — local changes to a small part of the model that explain changes in the environment. In other words, learning a causal model amounts to learning how the world behaves as well as how the world can change. Capturing the intuition that most knowledge about the world can be reused even in the face of environment changes, the Sparse Mechanism Shift hypothesis [39, 49, 3] posits that natural changes in the data distribution can be explained by sparse changes in the causal mechanisms. This suggests that a world model reflecting the causal structure of the world can adapt efficiently to changes by updating only a small part of the model while most of the model remains invariant.
In this light, several recent works have explored the benefits of causally structured world models [e.g. 33, 20]. These approaches aim to learn a causal graph that captures how entities in an environment influence each other. However, these causal discovery methods learn a fixed graph to explain the entire dataset, which suffer from two principle drawbacks in the context of modeling realistic physical systems. First, learning a fixed causal graph that explains the data requires the graph capture all possible interactions between entities in the scene. In most realistic settings, where all objects can potentially interact with each other, this often leads to uninformative, fully-connected graphs. Second, in many application domains, such as traffic motion prediction [52], the number of entities can change from scene to scene, which is incompatible with the standard structure learning approach. In reality, most physical interactions such as collisions between objects manifest in temporally sparse events. As such, we argue that local causal models [42, 43, 50, 21], which capture only the relevant causal relationships at each timestep, provide a more natural and flexible framework for learning structured world models. To this end, [42] propose that local causal graphs can be inferred by investigating the attention pattern that emerges in a Transformer-based dynamics model [56]. While this is sufficient in simple, state-based settings, we argue and empirically show that attention alone is not able to reliably discover local causal relationships in environments with more complex dynamics or high-dimensional observations. To remedy this, we take inspiration from optimisation-based causal discovery methods [5] which use sparsity to induce causal graphs, and use sparsity regularisation as an inductive bias to learn local causal models.
In this work, we develop a sparsity regularisation scheme that can be applied to the standard Transformer architecture, which enables the discovery of state-dependent and time-dependent local causal graphs. We present SPARTAN, a Transformer-based world model with learnable sparse local causal graph structure. Specifically, SPARTAN minimises the expected number of local causal edges between tokens, allowing the discovery of accurate and interpretable causal edges between objects. We evaluate our model on observation-based environments with physical interactions and a traffic motion prediction dataset [52], showing that SPARTAN is able to identify causal edges more accurately than prior approaches, resulting in significantly improved robustness against interventions on objects that are not causally related to the prediction. We further extend our model to explicitly represent causal interventions. In so doing, SPARTAN is able to efficiently adapt to changes in the environment in a few-shot manner. We demonstrate that due to the causal structure discovered, our model achieves better adaptation efficiency without compromising on overall prediction accuracy.
@@ 2 Background
We begin this section by introducing our problem setting and providing a brief outline of causal graphical models and causal discovery methods. These are then discussed in relation to our motivation in section 2.3, where we note in particular the role of causality in world models. The goal of our work is to learn a causal model that predicts future states from observation. In order to model interventions, we train our model on a set of intervened environments with unknown intervention targets. Specifically, the training data of our model consists of observation sequences sampled from different environments $\{\mathbf{x}^{0},\mathbf{x}^{1},...,\mathbf{x}^{T},I\}$, where $x^{t}$ is the observation at timestep $t$ and $I$ is the environment index. Here, each environment is obtained by altering the dynamics of the system such as changing the friction of an object. We assume access to a representation model that maps observations to latent states. We further assume that the latent state space, $\mathcal{S}$ can be factorised, $\mathcal{S}=\mathcal{S}_{i}\times...\times\mathcal{S}_{N}$, such that each $\mathcal{S}_{i}$ represents the state of an entity such as an object in the scene. Examples of such representation models include object-centric models [37, 13, 12, 60] or disentangled causal representations [49, 36, 29, 32, 10] where each $\mathcal{S}_{i}$ is an object slot or a causal variable respectively. Our aim is to learn a transition function $p(s^{t+1}|s^{t})$ that captures the local causal structure.
@@ 2.1 Causal Graphical Models
A causal graphical model (CGM) [41] consists of a set of random variables $\mathcal{V}=\{V_{1},...,V_{N}\}$, a directed acyclic graph $\mathcal{G}=(V,E)$ over the set of nodes $\mathcal{V}$ and a conditional distribution for each random variable $V_{i}$, $P(V_{i}|Pa(V_{i}))$, where $Pa(V_{i})$ is the set of parents of $V_{i}$. Each edge in the graph $(i,j)$ means that $V_{i}$ is a cause of $V_{j}$. The joint distribution over the variables can be factorised as
@@
where $P^{\prime}(\cdot)$ is the intervened conditional distribution. In contrast to a probabilistic graphical model, given a set of interventions, a CGM defines a family of distributions instead of a single, fixed one.
@@
2.2 Causal Discovery
The goal of our work is to discover local causal structures from observations. We take as starting point causal discovery methods, which aim to identify causal structures with data. We take inspiration from the continuous optimisation formulation [62, 5], which can be naturally incorporated into the framework of world models. Given a set of interventions with unknown targets $\mathcal{I}^{1:K}$ and their corresponding distributions $X\sim p_{X}^{(k)}$, causal discovery can be formulated as the following [5],
$\max_{\theta}\sum^{K}_{k=1}\mathbb{E}_{X\sim p_{X}^{(k)}}\log f^{(k)}(X;\theta)-\lambda|\mathcal{G}(\theta)|-\lambda_{R}|\mathcal{I}(\theta)|,$ (3)
where $\theta$ are the learnable parameters, $f^{(k)}(\cdot)$ is the likelihood of the data under the model for the $k$-th intervention, $|\mathcal{G}(\theta)|$ is the number of edges in the learnt causal graph, $|\mathcal{I}(\theta)|$ is the number of intervention targets, and $\lambda$ and $\lambda_{R}$ are hyperparameters for sparsity regularisation.
Intuitively, solving this optimisation amounts to pruning out spurious edges and interventions, and finding a model that can explain the data with the sparsest causal graph and intervention targets. Analogously, we apply sparsity regularisation in the setting of learning local causal models. As discussed above, local causal graphs and intervention targets are induced by the state. Here, the goal is to infer the sparsest graph at each timestep such that the next observation can be explained. To this end, we modify the optimisation objective for local causal discovery: instead of minimising the size of a fixed global causal graph and intervention targets, we propose to minimise the expected number of edges and targets with respect to the data distribution where the causal edges and intervention targets are built dynamically at each timestep. In Sec. 3, we present our main model, SPARTAN, which is an instantiation of this idea using a Transformer-based architecture.
@@ 2.3 Causality and World Models
Conceptually, we argue that in order to perform efficient adaptation, world models should be structured to reflect the underlying sparse causal structure of the observed dynamics, and that these structures should be local. To illustrate our point, consider the setting of traffic behaviours: changes in traffic rules arise when travelling to countries where vehicles drive on the other side of the road. Although the relative position of vehicles given the road boundaries may change, most traffic behaviours (e.g. stopping at a red light or lane keeping) would remain the same. A world model that can efficiently adapt in this setting should modularly update only a small subset of learnt dynamics (i.e. the relative position of vehicles with respect to the road boundary) while leaving other dynamics unchanged. Our proposition is motivated by the sparse mechanism shift hypothesis [39, 49, 3], which states that naturally occurring distribution changes can be attributed to sparse interventions on causal mechanisms. Formally, this means that most reasonable changes in the environment can be modelled by changing a small subset of the conditional distributions in Eq. 1 while the rest stay invariant. Furthermore, in the special case where interventions act only on variables that are not causally related to model predictions, e.g. removing irrelevant objects in a scene [47], a model that reflects the correct causal structure would remain robust to changes.
While learning a global causal graph is sufficient in some settings, in the context of dynamics models, a global graph is often close to fully connected since any entities that interact with each other at any point are connected on the graph regardless of how unlikely the interaction is. Returning to the traffic example, a global causal graph would connect every vehicle in the scene as they can all influence each other when close together. However, events such as "vehicle A causes vehicle B to stop" can be more appropriately captured by a local causal graph with the edge $A\rightarrow B$ but not edges to other vehicles that are irrelevant at the time. Here we posit that a global graph containing many locally irrelevant edges prevents the model from fully exploiting the sparse structure of the problem at hand. In the following sections, we present SPARTAN. We demonstrate experimentally that learning a sparse, local causal world model improves robustness and adaptation efficiency, and that our model is capable of discovering causal structures accurately.
@@ 3 Sparse Transformer World Models
Our goal is to develop a world model that learns local causal models as transition functions. We build on the Transformer architecture [56] as it achieves state-of-the-art prediction performance in object-factored world models [61]. Moreover, its attention mechanism provides a natural way to control the flow of information between object tokens. While prior work [42] argues that soft attention is a strong enough inductive bias for learning local causal graphs and proposes a thresholding heuristic to the attention patterns, we posit that this is insufficient beyond simple state-based scenarios and that appropriate sparsity regularisation and hard attention plays a crucial role in scaling this up to more complex, observation-based environments. To demonstrate with a simple example, consider a system with only one causal edge, $s^{t}_{i}\rightarrow s^{t+1}_{j}$ and many other irrelevant nodes. A soft-attention model can to predict $s^{t+1}_{j}$ as long as the attention value for $s^{t}_{i}$ is non-zero, i.e. information can flow from the $i$-th token to the $j$-th token. However, the model is free to have non-zero attention values over other tokens. In this case, there is no guarantee that the attention value on $s^{t}_{i}$ is the highest amongst all tokens, and therefore applying a threshold on the attention values may ’catch’ spurious edges or miss neccessary edges. Instead, a model that masks the information flow between the tokens and we penalise connections between tokens. Any model that contains spurious connections would have a high loss due to the sparsity penalty while models that do not contain the $s^{t}_{i}\rightarrow s^{t+1}_{j}$ would suffer from low prediction accuracy. Under this scheme, the model that achieves the best training loss would be the one that only contains the correct causal edge. In the following section, we present SPARrse TrANsformer World models (SPARTAN), an instantiation of a masked transition model using a Transformer-based architecture with hard attention.
@@ 3.1 Learnable Sparse Connections
We start with the architecture of a single layer Transformer before extending it to the multi-layer case. Given object representations $s^{t}_{1:N}$ as input, we obtain the keys, queries and values, $\{k_{i},q_{i},v_{i}\}$ via linear projection as in the standard Transformer. Using these, we sample a binary adjacency matrix,
@@
where $\text{Bern}(\cdot)$ is the Bernoulli distribution and $\sigma$ is the sigmoid function. The hidden features are then computed using the standard scaled dot-product attention with the the adjacency matrix as mask.
@@
where $\hat{s}^{t+1}_{i}$ is the prediction for the $i$-th object in the next timestep. Here, the adjacency matrix acts as mask that disallows information to flow from tokens that are not in the parent set of the querying token. Note that the adjacency matrix is dependent on the current state via the key-query pairs. The learnt adjacency matrix can therefore be interpreted as the local causal graph, i.e. $A_{ij}=1$ means $s^{t}_{j}\in Pa^{\mathcal{L}}(s^{t+1}_{i})$. The sampling step is made differentiable via Gumbel softmax trick [23]. Typically, multiple Transformer layers are stacked sequentially for expressivity. This presents a problem for the learnt adjacency matrix as information can flow between tokens in a multi-hop manner across different layers, e.g. token $i$ can influence token $j$ via $i\rightarrow k\rightarrow j$ without the edge $(i,j)$ being in any of the adjacency matrices associated with different layers. SPARTAN mitigates this problem by tracking the number of paths that leads from token $i$ to $j$. Concretely, for a $L$-layer Transformer, we compute the path matrix111[31, 42] use a similar formulation of path matrices for bounding the Jacobian between the tokens.,
@@
where $A^{l}$ is the adjacency matrix at layer $l$ and $\mathbb{I}$ is the identity matrix. Here the identity matrix is added because by default each token has a self edge due to residual connections in the Transformer architecture. The path matrix $\bar{A}$ has the property that $\bar{A}_{ij}$ is the number of paths that leads from $j$ to $i$. In this case, $s^{t}_{j}$ is a local causal parent of $s^{t+1}_{i}$ iff $\bar{A}_{ij}>=1$.

### v3

# SPARTAN: A Sparse Transformer World Model Attending to What Matters

@@ Abstract
Capturing the interactions between entities in a structured way plays a central role in world models that flexibly adapt to changes in the environment. Recent works motivate the benefits of models that explicitly represent the structure of interactions and formulate the problem as discovering local causal structures. In this work, we demonstrate that reliably capturing these relationships in complex settings remains challenging. To remedy this shortcoming, we postulate that sparsity is a critical ingredient for the discovery of such local structures. To this end, we present the SPARse TrANsformer World model (SPARTAN), a Transformer-based world model that learns context-dependent interaction structures between entities in a scene. By applying sparsity regularisation on the attention patterns between object-factored tokens, SPARTAN learns sparse, context-dependent interaction graphs that accurately predict future object states. We further extend our model to adapt to sparse interventions with unknown targets in the dynamics of the environment. This results in a highly interpretable world model that can efficiently adapt to changes. Empirically, we evaluate SPARTAN against the current state-of-the-art in object-centric world models in observation-based environments and demonstrate that our model can learn local causal graphs that accurately reflect the underlying interactions between objects, achieving significantly improved few-shot adaptation to dynamics changes, as well as robustness against distractors.
@@ 1 Introduction
World Models [ha2018recurrent] have emerged in recent years as a promising paradigm to enable a wide range of downstream tasks such as video prediction [villegas2019high, gaia], physical reasoning [chen2021grounding], and model-based RL [hafner2020dreamerv2]. Recent advances have employed the transformer architecture [attention_is_all_you_need] to develop world models capable of performing accurate predictions over ever longer horizons in increasingly complex settings [micheli2023transformers, robine2023transformerbased]. Owing to the flexibility of attention mechanisms, the transformer architecture can be combined with object-centric representations [slot_attention] to accurately capture interactions between objects, achieving state-of-the-art prediction performance [wu2023slotformer]. However, the ability to generalise and adapt to changes in the environment in a data-efficient manner remains a significant challenge. Evidence suggests that transformer world models can be sensitive to changes in distractors [causal_agents], further highlighting the need for models to capture robust and generalisable interactions. As such, recent works have argued that explicitly representing interaction structures between entities, i.e. when and how entities influence each other, is critical in improving generalisation [pitis2022mocoda] and adaptation efficiency [lei2023variational, huang2021adarl].
To this end, the intersection between causality and machine learning [scholkopf2021toward] offers a promising framework for building structured models that capture interactions in a principled manner. Through this lens, the forward dynamics can be factorised into independent causal mechanisms, where prediction for each entity is made based solely on other entities that have a causal influence on it. Here, the Sparse Mechanism Shift hypothesis [1911.10500, Bengio2020A] posits that natural changes in the data distribution can be explained by sparse changes in these causal mechanisms. This suggests that a world model reflecting the causal structure of the world can adapt efficiently to changes by updating only a small part of the model while most of the model remains invariant. This causal view of world modelling has motivated prior works which explored the benefits of structured world models [lei2023variational, huang2021adarl]. These approaches aim to learn a graph that encodes how entities in an environment influence each other. However, these methods learn a fixed graph to fit an entire dataset, which is unsuitable for modelling realistic physical interactions, such as collisions between objects, that manifest as temporally sparse events. To remedy this, recent approaches [seitzer2021causal, hwang2024finegrained] formulate the problem as learning state-dependent local causal graphs (Fig. 1). Despite their success in simple state-based settings, scaling prior local causal discovery methods to image observations with complex dynamics remains a significant bottleneck. This hinders the application of these methods to realistic scenarios where, for example, objects are represented via learned embeddings and the number of objects can vary from scene to scene.
Taking inspiration from the conceptual framework of local causal models, the goal of this work is to develop a transformer-based world model that faithfully reflects the interaction structure between objects while maintaining the flexibility and accuracy afforded by the transformer architecture. The key insight is that attention mechanisms serve as a natural starting point for constructing state-dependent graphs. By applying sparsity regularisation, an inductive bias used in prior local causal discovery methods [hwang2024finegrained], on a modified transformer architecture with hard attention, we develop SPARTAN, a Transformer-based world model with a learnable local causal graph. We evaluate our model on observation-based environments with physical interactions and a real-life traffic motion prediction dataset [waymo]. To the best of our knowledge, SPARTAN is the first method that learns local causal structures beyond simple state-based settings. Compared to prior transformer world models, our results show that SPARTAN is able to accurately identify causal edges, resulting in significantly improved robustness against distractors, few-shot adaptation, and interpretability.
@@ 2 Background
We begin this section by introducing our problem setting and providing a brief outline of causal graphical models and causal discovery methods. These are then discussed in relation to our motivation in section 2.3, where we note in particular the role of causality in world models. The goal of our work is to learn a causal model that predicts future states from observation. In order to model interventions, we train our model on a set of intervened environments with unknown intervention targets. Specifically, the training data of our model consists of observation sequences sampled from different environments $\{\mathbf{x}^{0},\mathbf{x}^{1},...,\mathbf{x}^{T},I\}$, where $x^{t}$ is the observation at timestep $t$ and $I$ is the environment index. Here, an intervention amounts to altering the dynamics such as changing the friction of an object. We assume access to a representation model that maps observations to factored latent states. $\mathcal{S}=\mathcal{S}_{i}\times...\times\mathcal{S}_{N}$, such that each $\mathcal{S}_{i}$ is an embedding that represents the state of an entity such as an object in the scene. Examples of such representation models include object-centric models [slot_attention, engelcke2021genesisv2] or disentangled causal representations [1911.10500, lippe2022citris, Leebetal23] where each $\mathcal{S}_{i}$ is an object slot or a causal variable respectively. Our aim is to learn a transition function $p(s^{t+1}|s^{t})$ that captures the local causal structure.
@@ 2.1 Causal Graphical Models
A causal graphical model (CGM) [PetJanSch17] consists of a set of random variables $\mathcal{V}=\{V_{1},...,V_{N}\}$, a directed acyclic graph $\mathcal{G}=(V,E)$ over the set of nodes $\mathcal{V}$ and a conditional distribution for each random variable $V_{i}$, $P(V_{i}|Pa(V_{i}))$, where $Pa(V_{i})$ is the set of parents of $V_{i}$. Each edge in the graph $(i,j)$ means that $V_{i}$ is a cause of $V_{j}$. The joint distribution over the variables can be factorised as
@@
where $P^{\prime}(\cdot)$ is the intervened conditional distribution. In contrast to a probabilistic graphical model, given a set of interventions, a CGM defines a family of distributions.
Recent works in causal discovery, which aim to learn causal graph structure from data, formulate the identification of causal structures as sparsity-regularised optimisation [zheng2018dags, brouillard2020differentiable]. Intuitively, this amounts to pruning spurious edges and interventions and finding a model that can explain the data with the sparsest causal graph. Building on these works, we use sparsity regularisation as an inductive bias to learn local causal graphs.
2.2 Local Causal Models
@@
Analogously to standard causal discovery, our goal is to infer the sparsest graph at each timestep, such that the next observation can be explained. In the local case, the optimisation objective for local causal discovery is to minimise the expected number of edges and targets with respect to the data distribution, where the causal edges and intervention targets are state-dependent and are built dynamically at each timestep. [hwang2024finegrained] provides a theoretical analysis of the conditions under which local causal structures can be identified using sparse regularisation.
@@ 2.3 Causality and World Models
Conceptually, we argue that world models should be structured to reflect the underlying sparse and local causal structure of the dynamics to perform efficient adaptation. To illustrate our point, consider the setting of traffic behaviours: changes in traffic rules arise when travelling to countries where vehicles drive on the opposite side of the road. Although the relative position of vehicles given the road boundaries may change, most traffic behaviours (e.g., stopping at a red light or lane keeping) would remain the same. A world model that can efficiently adapt should update only a small subset of learnt dynamics (i.e., the relative position of vehicles to the road boundary) while leaving other dynamics unchanged. Our proposition is motivated by the sparse mechanism shift hypothesis [ParKilRojSch18, 1911.10500], which states that naturally occurring distribution shifts can be attributed to sparse interventions on causal mechanisms. Formally, this means that reasonable changes in the environment can be modelled by changing a small subset of the conditional distributions in Eq. 1 while the rest remain invariant. In the special case where interventions act only on variables that are not causally related to model predictions, e.g., removing irrelevant objects in a scene [causal_agents], models that reflect the correct causal structure would remain robust.
While learning a global causal graph is sufficient in some settings, in the context of dynamics models, such a graph is often close to fully connected since any entities that interact with each other at any point are connected in the graph regardless of how unlikely the interaction is. Returning to the traffic example, a global causal graph would connect every vehicle in the scene, as they can all influence each other when close together. However, events such as "vehicle A causes vehicle B to stop" can be more appropriately captured by local causal graphs with the edge $A\rightarrow B$ but not by edges to other vehicles that are irrelevant at the time. As such, we argue that local causal graphs are more suitable for models to fully exploit the sparse structure of the problem at hand. In the following section, we present SPARTAN, a transformer-based instantiation of sparsity-regularised local causal discovery.
@@ 3 Sparse Transformer World Models
Our goal is to develop a world model that learns local causal models as transition functions. We build on the Transformer architecture [attention_is_all_you_need] as it achieves state-of-the-art prediction performance in object-factored world models [wu2023slotformer]. Moreover, its attention mechanism provides a natural way to control the flow of information between object tokens. While prior work [pitis2020count] argues that soft attention is a strong enough inductive bias for learning local causal graphs and proposes a thresholding heuristic on the attention patterns, we posit that this is insufficient beyond simple state-based scenarios and that appropriate sparsity regularisation and hard attention play a crucial role in scaling this up to more complex, observation-based environments. To demonstrate with a simple example, consider a system with only one causal edge, $s^{t}_{i}\rightarrow s^{t+1}_{j}$ and many other irrelevant nodes. A soft-attention model can predict $s^{t+1}_{j}$ as long as the attention value for $s^{t}_{i}$ is non-zero, i.e., information can flow from the $i$-th token to the $j$-th token. However, the model is free to have non-zero attention values over other tokens. In this case, there is no guarantee that the attention value on $s^{t}_{i}$ is the highest amongst all tokens, and therefore, applying a threshold on the attention values may ’catch’ spurious edges or miss necessary edges. Instead, we need a model that masks the information flow between the tokens and penalises connections between tokens. Any model that contains spurious connections would have a high loss due to the sparsity penalty, while models that do not contain the $s^{t}_{i}\rightarrow s^{t+1}_{j}$ would suffer from low prediction accuracy. Under this scheme, the model that achieves the best training loss would be the one that contains only the correct causal edge. In the following section, we present SPARrse TrANsformer World models (SPARTAN), an instantiation of a masked transition model using a Transformer-based architecture with hard attention.
@@ 3.1 Learnable Sparse Connections
We start with the architecture of a single layer before extending the apporach to the multi-layer case. Given object representations $s^{t}_{1:N}$ as input, we obtain the keys, queries, and values, $\{k_{i},q_{i},v_{i}\}$ via linear projection, as in the standard Transformer. Using these, we sample an adjacency matrix,
@@
where $\text{Bern}(\cdot)$ is the Bernoulli distribution and $\sigma$ is the sigmoid function. The hidden features are then computed using the standard scaled dot-product attention with the adjacency matrix as masks.
@@
where $\hat{s}^{t+1}_{i}$ is the prediction for the $i$-th object in the next timestep. Here, the adjacency matrix acts as a mask that disallows information flows from tokens that are not in the parent set of the querying token. Note that the adjacency matrix is dependent on the current state via the key-query pairs. The learnt adjacency matrix can therefore be interpreted as the local causal graph, i.e. $A_{ij}=1$ means $s^{t}_{j}\in Pa^{\mathcal{L}}(s^{t+1}_{i})$. The sampling step is made differentiable via Gumbel softmax trick [Gumbel_max]. Typically, multiple Transformer layers are stacked sequentially for expressiveness. This presents a problem for the learnt adjacency matrix, as information can flow between tokens in a multi-hop manner across different layers: for example, token $i$ can influence token $j$ via $i\rightarrow k\rightarrow j$ without the edge $(i,j)$ being present in any of the adjacency matrices. SPARTAN mitigates this problem by tracking the number of paths from token $i$ to $j$. Concretely, for $L$ layers, we compute the path matrix,
@@
where $A^{l}$ is the adjacency matrix at layer $l$ and $\mathbb{I}$ is the identity matrix. The identity matrix is added due to residual connections. The path matrix $\bar{A}$ has the property that $\bar{A}_{ij}$ is the number of paths that lead from $j$ to $i$. In this case, $s^{t}_{j}$ is a local causal parent of $s^{t+1}_{i}$ iff $\bar{A}_{ij}>=1$.